two trains travel in opposite directions from the same station at rates differing by 20 kph. in two hours they will be 160 km apart
Answer by mananth(16949)

Train A speed =x
Train B speed = x+20
they are moving away
so add up speeds
combined speed = 2x+20
distance = 160
time = 2hours
t=d/r
2=160/(2x+20)
2(2x+20)=160
4x+40=160
4x=120
/4
x=30 mph
The speeds are 30 & 50 mph